OpenStudy (anonymous):

I'm confused how to find the least common multiple of x^2-4x, x^2+4x. I know they factor out to x(x-4) and x(x+4) but dont know how to go any further

OpenStudy (anonymous):

do you know how to find the gcd of the two functions?

OpenStudy (anonymous):

the gcd is the terms that are common between the two functions

OpenStudy (anonymous):

in this case, x

OpenStudy (anonymous):

Yes, i got that far.

OpenStudy (anonymous):

to find the lcm, multiple the functions together and divide by the gcd

OpenStudy (anonymous):

*multiply

OpenStudy (anonymous):

so the lcm would be x(x-4)x(x+4)/x, which is equal to x(x-4)(x+4)

OpenStudy (anonymous):

ok that makes no sense lol

OpenStudy (anonymous):

Maybe you can clarify it a little better for me, looking at a problem x^2-6x, x^2 i know the lcm is x^2(x-15) how is this? i know this is correct bc of my text.

OpenStudy (anonymous):

i think your text has a typo. the answer should be x(x^2-16)

OpenStudy (anonymous):

hmmmmm

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

lcm of x^2-6x and x^2 should be x^2(x-6) x^2-6x = x(x-6) x^2 = x.x gcd = x carry out the method suggensted by eashy : LCM = x(x-6)*x^2/x = x^2(x-6)

OpenStudy (amistre64):

one idea is to spose you are try to add fractions, with those as denominators: \[\frac{1}{x^2-4x}+\frac{1}{x^2+4x}=n\] factor them \[\frac{1}{x(x-4)}+\frac{1}{x(x+4)}=n\] start clearing out the fractions by multiplying thru by something useful, like an x, and simplifying the results \[\frac{\cancel x}{\cancel x~(x-4)}+\frac{\cancel x}{\cancel x~(x+4)}=x~n\] \[\frac{\cancel{(x-4)}}{\cancel{(x-4)}}+\frac{(x-4)}{(x+4)}=x(x-4)~n\] \[\frac{(x-4)\cancel{(x+4)}}{\cancel{(x+4)}}=x(x-4)(x+4)~n\] the LCM gets created on the right hand side next to that "n" once you have cleared all the fractions